# Isentropic processes for axisymmetric Black Holes

### v1

@@ 4 Discussion
We investigated whether a stationary axisymmetric black hole can absorb a classical test charged particle without a change in the black hole’s entropy, and found that this process is classically forbidden in any diffeomorphism-invariant theory of gravity. Further, we analysed a Kerr-Newman blackhole as a specific example. A test charged particle outside a Kerr–Newman black hole experiences an effective potential for isentropic absorption, which remains positive between the outer horizon and a certain distance away from it. So, classically, this region outside the black hole is forbidden. However, for a certain range of mass, charge, and angular momentum of a test charged particle, it is quantum mechanically possible to tunnel through this classically forbidden region.
@@
Our results show that the tunneling probability increases with a decrease in mass of either the test particle or the black hole. However, the tunneling probability increases with an increase in the charge of either the test particle or the black hole. Furthermore, we also find that the tunneling probability is large for a small blackhole in comparison to the larger black holes. These observations suggest that the low mass and high charge black holes with a smaller radius of the outer horizon are more prone to isentropic absorption of classical test charged particles. Moreover, if a charged particle that is entangled with degrees of freedom outside the black hole is isentropically thrown into the black hole, the entanglement between the inside and outside of the black hole increases without any change in the areal entropy. As emphasized in [28], this raises important questions about whether the black hole can accumulate entanglement entropy isentropically over time. Since the Boltzmann entropy accounts for the total number of accessible microstates and is typically greater than the entanglement entropy [28, 43, 44, 45], could it perhaps happen through such a process that the Boltzmann entropy becomes greater than the Bekenstein-Hawking entropy?

### v2

@@ 4 Discussion
We investigated whether a stationary axisymmetric black hole can absorb a classical test charged particle without a change in the black hole’s entropy, and found that this process is classically forbidden in any diffeomorphism-invariant theory of gravity. We found that the conditions required to keep the Bekenstein–Hawking entropy unchanged lead to a classically inaccessible region outside the Killing horizon. Further, we analysed a Kerr-Newman black hole as a specific example. A test charged particle outside a Kerr–Newman black hole experiences an effective potential for isentropic absorption, which remains positive between the outer horizon and a certain distance away from it. So, classically, this region outside the black hole is forbidden. However, for a certain range of mass, charge, and angular momentum of a test charged particle, it is quantum mechanically possible to tunnel through this classically forbidden region.
@@
Our results show that the tunnelling probability increases with a decrease in the mass of either the test particle or the black hole. However, the tunnelling probability increases with an increase in the charge of either the test particle or the black hole. Furthermore, we also find that the tunnelling probability is large for a small black hole in comparison to the larger black holes. These observations suggest that the low mass and high charge black holes with a smaller radius of the outer horizon are more prone to isentropic absorption of classical test charged particles. Moreover, if a charged particle that is entangled with degrees of freedom outside the black hole is isentropically thrown into the black hole, the entanglement between the inside and outside of the black hole increases without any change in the areal entropy. As emphasised in [28], this raises important questions about whether the black hole can accumulate entanglement entropy isentropically over time. Since the Boltzmann entropy accounts for the total number of accessible microstates and is typically greater than the entanglement entropy [28, 44, 45, 46], could it perhaps happen through such a process that the Boltzmann entropy becomes greater than the Bekenstein-Hawking entropy?
From the perspective of information theory, the existence of such channels suggests that the information content associated with black holes may not always be constrained in the strict semiclassical sense. This can be contrasted with earlier work linking entropy bounds to information capacity [5, 22] and to the redistribution of fundamental degrees of freedom [21]. Taken together, our results reinforce the view that the entropy-bound relation remains robust at the semiclassical level but may admit controlled, non-perturbative violations in black hole systems. Such violations are not pathological but instead highlight the limitations of semiclassical reasoning when extended over a black hole’s lifetime. They may also play a role in addressing the long-standing black hole information paradox [30, 27].
We discussed the isentropic absorption of charged particles by axisymmetric, non-extremal black holes in (3+1) dimensions. How our results change for extremal black holes remains a topic for future investigation, as this scenario may reveal new features of black hole thermodynamics and quantum gravity. Extending this analysis to rotating or charged black holes in higher dimensions may also uncover rich dynamics that are inaccessible in simpler (3+1)-dimensional models. We employed the saddle-point approximation of path integrals to evaluate the tunnelling probability, which provides a framework to study the leading-order contribution to the particle absorption beyond classical absorption. One could further explore the effects of loop corrections, as well as backreaction effects, on the tunnelling probabilities, potentially leading to refined predictions for black hole evaporation rates. Another direction is to investigate whether these non-perturbative channels leave observable imprints in late-stage Hawking radiation or in quantum-gravitational corrections to entropy bounds, which could provide novel tests for our understanding of information loss in black holes.